For a given motion, the relationship between time t and distance x is found out to be t=ax^2+bx, where a and b are constants. Considering v as velocity, the retardation will be given by,
1
Expert's answer
2019-06-03T10:17:02-0400
"t=ax^2+bx"

Differentiate both side with respect to time:

"1=2ax\frac{dx}{dt}+ b\frac{dx}{dt}=2axv+bv"

"v^{-1}=2ax+b"

Differentiate both side with respect to time:


"-v^{-2}\frac{dv}{dt}=2a\frac{dx}{dt}=2av"

Th retardation is


"-\frac{dv}{dt}=2av^3"